an engine can pull 4 coaches at a maximum speed of 20m/s. mass of the engine is twice the mass of every coach. Assuming resistive forces to be proportional to the weight, approximate maximum speed of the engine when it pulls 12 and 6 coaches are what ?

Grade:11

3 Answers

Vikas TU
7 years ago
From WET,
(4N + 2N)x = 0.5mv^2
6mgx = mv^2/2 {Resistive force is prop. to  weight}
x = 10/3 meter.
Now for 12 coaches,
(12 + 2)mgx = 0.5mv^2
we get,
v = 30.123 m/s (approx.)
for 6 coaches,
8mgx = 0.5mv^2
v = 23.1245 m/s (approx.)
Harsh
6 years ago
Total mass; f is directly proportional to 6m, f=6 (m of coach)20=PTo drive 12m:f is directly proportional to 14m =>f=14 (mass of coach) 14×mass of coach×v = 6×mass of coach×20 =>v=8.57m/sSimilarly to drive 6 boggies: f is directly proportional to 8mSimilarly calculate P and then v=15m/s
Nilabh
5 years ago
A/Q 
  P=k 6mgv----------------(1)
  P=k 14mgv1----------------(2)
From eqn 1 and 2:-
 
6mgv=14mgv1
6v=14v1
V1=6×20/14
V1=8.5m/s
 
 
Also,
6mgv=8mgv2
6v=8v2
V2=6×20/8
V2=15m/s